Die Korrespondenz Helmut Hasse/Constance Reid

tk Hasse an Reid 10.10.66–5.6.76
tk Reid an Hasse 10.8.66–24.5.76
tk Weiteres Material

t – fertig transkribiert, k – nach Tippfehlern durchgesehen

Version vom 26.4.2007

Letztmalig geändert am 29.4.2007

Quelltext: hascre_070426.tex übersetzt am 21. Juli 2014

Inhaltsverzeichnis

1 Die Korrespondenz Helmut Hasse/Constance Reid 5

1.1 10.08.1966, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 6

Asking for information about Hilbert’s life after 1933.

1.2 10.10.1966, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 7

Some stories about Hilbert. Hilbert on Königsberg. H. explaining to Hilbert the new face of class field theory. Hilbert meets young mathematicians on H.s group theory conference 1939.

1.3 23.10.1966, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 11

Thanks for reply. R. explains her concept of Hilbert book.

1.4 02.11.1966, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 12

Another Hilbert story. H. met Hilbert first in Leipzig 1922. H.s personal reaction to Hilbert’s work.

1.5 10.01.1968, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 14

H. now back in Germany. R. has finished the manuscript on Hilbert. Where are Hilbert’s letters to Minkowski? Does H. know about the circumstances of Hilbert’s name appearing on the proclamation in favor of Hitler?

1.6 28.01.1968, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 16

H. asked Grunwald about Hilbert’s letters to Minkowski. About the 1934 proclamation.

1.7 21.01.1970, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 17

R. sends H. her book on Hilbert, with many thanks.

1.8 01.02.1970, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 18

Thanks for the book. Some comments. About Hilbert’s lecture course 1919 on “Anschauliche Geometry”. H. explains his political attitude. H. had not been a “nationalist” but considered himself as a national citizen. H. had never been a member of any political party.

1.9 20.02.1970, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 20

R. has already a file of corrections and would appreciate receiving addtional remarks. The word “nationalist” is not meant in a derogatory sense.

1.10 29.12.1970, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 21

H. is in San Diego. R. would like to meet H. and talk further about Hilbert in his last years.

1.11 18.03.1971, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 22

Trying to find a date for meeting Hasse. H. proposes to meet on 24 April at Fibonacci conference in San Francisco.

1.12 28.03.1971, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 23

Will meet H. on 24 April in S.F.

1.13 17.04.1976, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 24

R. finished her book on Courant. R. has asked Springer to send galley proofs to H., so he would be able to see how R. treats H.’s role in Göttingen after 1933.

1.14 09.05.1976, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 25

H. thanks for Galley proofs. Looking forward to have full story of Courant’s life.

1.15 11.05.1976, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 26

H. has received more galley proofs. H. points out that there are statements which are not correct and asks for correction. About H.’s application for membership in the Nazi party; this was never granted. About H.’s activities in Göttingen. H. could help mathematicians who had problems of political nature. About the happenings in Göttingen in 1945/46.

1.16 12.05.1976, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 28

Additional remarks to yesterday’s letter. H. explains his political feelings. He has never been a member of any (political) Party. H. explains his reaction at the Faculty meeting in 1945. H. was strongly opposed to the antisemitism of Hitler and his party. H. names some jewish friends, among them Hensel, Courant and Toeplitz.

1.17 19.05.1976, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 30

R. accepts H.s statement regarding his status as party member. R. encloses a statement of the Berlin Document Center on that point. R. has tried to present controversial material—which comes up frequently in connection with Courant himself—from the point of view of both sides.

1.18 24.05.1976, Reid an Hasse . . . . . . . . . . . . . . . . . . . . 32

R. encloses the revisions in her galleys on the basis of H.s letters. Also another paragraph where H. is mentioned in relation to Artin.

1.19 05.06.1976, Hasse an Reid . . . . . . . . . . . . . . . . . . . . 33

H. replies and details certain controversial material in her book.

2 Weiteres Material zu Hasse/Reid 35 2.1 18.08.1934, Proklamation . . . . . . . . . . . . . . . . . . . . . 36 2.2 13.04.1972, Berlin Docum. Center an Reid . . . . . . . . . . . 38 2.3 03.05.1976, Springer an Hasse . . . . . . . . . . . . . . . . . . 39 2.4 05.05.1976, Springer an Hasse . . . . . . . . . . . . . . . . . . 40 2.5 07.05.1976, Springer an Hasse . . . . . . . . . . . . . . . . . . 41

3 Register 42

Kapitel 1 Die Korrespondenz Helmut Hasse/Constance Reid

\underline{10.08.1966, Reid an Hasse} 6

1.1 10.08.1966, Reid an Hasse

Asking for information about Hilbert’s life after 1933.

70 Piedmont Street, San Francisco, Calif., USA

August 10, 1966

Dear Professor Hasse:^{1}

In spite of David Hilbert’s great importance for twentieth century mathematics, there is almost no personal or biographical material about him available in English. For this reason, I have written a short, essentially non–technical biography which I hope will re–create for students and the general public the human–scientific phenomenon of Hilbert and of the Mathematical Institute in Göttingen. Both Professor Courant and Professor Bernays have offered to read my manuscript to make sure that the flavor is authentic. I have not been able to obtain much information about Hilbert’s life after 1933. For that reason I would be especially interested to know about your own association with Hilbert, your own direct personal impression of him, and any incidents which you feel give a picture of the man and his mind. May I express my deep appreciation for any help which you feel that you can give me in this matter?

Sincerely yours,

Constance Reid

Mrs. Neil D. Reid

1Vermerk von Hasse: Legh Wilber Reid, Dissertation 1899, Tafeln d. Klassenzahlen kub. Zahlkörper

\underline{10.10.1966, Hasse an Reid} 7

1.2 10.10.1966, Hasse an Reid

Some stories about Hilbert. Hilbert on Königsberg. H. explaining to Hilbert the new face of class field theory. Hilbert meets young mathematicians on H.s group theory conference 1939.

October 10, 1966

Dear Mrs. Reid:

Excuse me for having delayed answering your letter of August 10^{▶} until now. The preparations for my trip here where I spend the running academic year as Visiting Professor—and the trip itself as well as settling down here made me neglect my correspondence.^{1} I suppose you are a relative of Dr. Legh Wilber Reid whose Dissertation of 1899 “Tafeln der Klassenzahlen kubischer Zahlkoerper”, written under Hilbert’s supervision, is on my bookshelf since the early twenties when I began studying mathematics and found the volume in a Goettingen bookshop. You want to have information about Hilbert’s life after 1933. Although I came to Goettingen in 1934 and stayed there as professor at the Mathematical Institute, I have not had much contact with Hilbert. He never came to the Institute in my time. He led a rather retired life with his wife and son and the domestic who, as far as I know lives in his house still today and takes care of his son whose mental faculties are rather restricted. I remember a few occasions where a number of mathematicians from Goettingen and other German universities gathered in Hilbert’s home to celebrate his birthday, so 1932 the seventieth and 1937 the seventyfifth. On the latter occasion he had invited also the two young nurses who treated him regularly with some kind of massage. He was sitting between them, his arms round their necks, while in the adjoining room somebody made a congratulatory speech. Hecke reminded him that he ought to listen to that speech, instead of talking to those girls, but he said: “Ach was, das ist v i e l schoener!” On another occasion of the same kind the conversation came to the question which German town was most beautiful. Some were in favour of Dresden, some others of Munich or further cities. Hilbert said at last: “Aber Nein, aber nein; die schoenste deutsche Stadt ist doch Koenigsberg”.. When his wife protested with: “Aber Davidchen, das kannst Du nun wirklich nicht sagen, Koenigsberg ist doch gar nicht so sehr schoen”, he replied fully convinced that he knew it best: “Aber Kaethe, i c h muss das doch wissen; ich hab’ doch mein ganzes Leben dort zugebracht”. And when his wife reminded him that they actually came to Goettingen already fourty years ago, he said: “Ach, die paar Jaehrchen! Mein ganzes Leben war ich in Koenigsberg!”. So his mind condensed all the fourty fruitful years of his wonderful achievements in so many branches of mathematics into a “paar Jaehrchen”, whereas his youth in Koenigsberg stuck out as a very long period. As you surely know, there are lots of anecdotes of this type about Hilbert, most of them true, some of them well invented. Those two I told you happened in my presence. I could tell you another one which happened in my presence, but I suppose you rather want more serious information about Hilbert’s last years. Unfortunately I cannot give you much of that. I know Hilbert went on trying to complete his proof that mathematical analysis is non–contradictory. In spite of his being retired, the Goettingen Institute had detached one of the assistants entirely at his disposal, viz., Dr. Arnold Schmidt. He regularly went to Hilbert’s home and worked with him on that question. There has not come much of these endeavours, though. If you have not already done so, you would better contact Dr. Arnold Schmidt, professor at the University of Marburg now. The only scientific contact I had with Hilbert was in the early thirties, I think in 1932. I had expressed to Mrs. Hilbert my ardent desire to talk once in my life personally to the great man. She was so kind as to arrange an opportunity for that by inviting me to tea and promising to leave me alone with him afterwards. So it happened. After tea he took me to his garden und we strolled to and fro underneath the green leaves of a wonderful pergola which, I believe, he had grown and cared for with his own hands in the course of the years. After having duly admired this beautiful garden, I began talking to him about what interested me most in those days, viz., the theory of algebraic numbers and in particular class–field theory. On this theory I had written a report, in continuation of Hilbert’s celebrated “Zahlbericht”, and I began telling him what I had done in this theory, based on his own famous results in the late nineties. But he interrupted me repeatedly and insisted that I explain him the basic conceptions and results of that theory before he could listen to what I wanted to tell him. So I explained to him the very foundations of class–field theory. About this he got extremely enthusiastic and said: “Das ist ja aber alles wunderschoen, wer hat denn das eigentlich gemacht ? ”. And I had to tell him that it was he himself who had laid that foundation and envisaged that beautiful theory. After that he listened to what I had to tell him about my own results. He listened attentively, more politely than intelligently. The conversation came then soon to an end, I think because he was called in by his wife because of some trivial matter.^{2} Just before the war, in July 1939 we had a colloquium about group theory at the Goettingen Institute.^{3} Quite a few experts on that theory from German and other European universities assisted, all rather young people. They asked me to procure them an opportunity to meet Hilbert and talk to him. So my wife and I invited Hilbert and his wife to our home to quite an informal dinner and a glass of wine or beer afterwards, and we invited also all the participants of our colloquium, altogether about 32 persons. We sat down in our two rooms downstairs and in my study upstairs. Whereas Mrs. Hilbert sat down in one of the downstairs rooms together with my wife and others, I directed Hilbert upstairs with a group of four or five of our other guests. So these latter had the opportunity of sitting close to him and have a good conversation with him. In the further course of the evening four or five others took their turn in the room upstairs, until everyone had had the desired opportunity. Hilbert was extremely gay and alert that evening. He liked young people far more than elderly ones. Although the conversation was mostly on general subjects other than maths, they all enjoyed this wonderful opportunity the memory of which will last for the rest of their lives. It was for all of us the last time we saw Hilbert. Six weeks later the war broke out which separated so many friends, families and nations for an indefinite time, and two and a half years later Hilbert died, having broken his arm on the street and never recovering from being constrained to lie in bed for several weeks, without his accustomed daily exercise.^{4} \underline{10.10.1966, Hasse an Reid} 10 I hope the few things I could tell you will be helpful to your purpose

1Hasse writes this letter from Honolulu.

2This anecdote is told in various versions, with various persons claiming to have been present. Here we learn, first that it really happened, and secondly that it was Hasse and not somebody else who had experienced this.

3The theme of the conference was p-groups, and Philip Hall was the main speaker. Hall delivered four lectures which were then published in Crelle’s Journal, volume 182.

4Hasse had been drafted to the Navy during the war. This may perhaps explain that he was not present at Hilbert’s funeral. But this may also be due to the fact that Hasse, as said above, had not much personal contact with Hilbert, so he had not been informed about the date and time of the funeral.

Yours very sincerely, H. Hasse

1.3 23.10.1966, Reid an Hasse

Thanks for reply. R. explains her concept of Hilbert book.

70 Piedmont Street, San Francisco, Calif. 94117

October 23, 1966

Dear Professor Hasse:

I hope I can express to you how delighted I was the other morning when I received your letter^{▶} and read the very poignant stories about Hilbert’s last years, which you recounted so well. I would be very much interested in hearing also the other anecdote which you mentioned but did not recount. So much of the man comes through in the authentic Hilbert stories that I want to keep a record of as many as possible, even though I shall not include all of them in my book. (I am very aware of the danger of reducing a man like Hilbert to simply a series of anecdotes.) As I mentioned to you in my first letter^{▶}, my book is to be a nontechnical portrait of Hilbert’s “mathematical personality”. In my view this includes both the personality of his mathematical work as well as the effect of his personality on other mathematicians and on mathematics during his time. For this reason I hope you will be generous enough to write again and tell me in as untechnical a way as possible your own personal reaction to Hilbert’s number theory work—a little of which I have already glimpsed in your letter. I would like my readers to glimpse something of the unique power and beauty of Hilbert’s mathematics, and it seems to me that this can come most forcibly through the firsthand reaction of someone like yourself who has actually built upon it.

May I thank you again for the assistance you have already given me.

Sincerely yours,

Constance Reid

P.S. Unfortunately I am no relation to Legh Wilber Reid, although I am familiar with his book on number theory and the very charming introduction which Hilbert wrote for it. I remember particularly his comment to the effect that every devotee of the theory desires it to be “equally a possession of all nations”.

1.4 02.11.1966, Hasse an Reid

Another Hilbert story. H. met Hilbert first in Leipzig 1922. H.s personal reaction to Hilbert’s work.

Nov.2, 1966

Dear Mrs. Reid:

The other story I had in mind also happened on the occasion of his birthday, I believe it was the 75 th. Somehow the conversation turned to the radio. Now on a previous birthday in the late twenties some big radio business firm had made him a present of their latest and promised to exchange it during the rest of his life always against their most recent radio. On the occasion mentioned Hilbert began to praise this wonderful invention of the human mind in ample emphatic words. He said, even the poorest family in Germany possessed a radio and listened in regularly all day, when on walking through the streets of Goettingen one found ever so many windows open and radios sounding at their loudest from within, at such a time Hilbert said: “Most people have not the slightest idea what a wonderful thing a radio is, they could not afford it, otherwise everybody would have one. As to myself, I am, as you know, in the enviable position of having installed in my home always the newest model of these precious valuable things. So I am really one of the very few happy persons who have the privilege to enjoy listening to music from a radio.”

Whereas this story—or rather these remarks—of Hilbert only show how far from actual life Hilbert’s mind was, at least in his last years, I was witness to another remark of Hilbert’s, made in 1922, which gives one a glimpse on his categorial attitude towards logic. It shows that he took this attitude not only when facing mathematical problems, where it was quite natural, but also in the field of human relations.

In September 1922 the Gesellschaft Deutscher Naturforscher und Aerzte held its first congress after the world war at Leipzig, and it was there that Hilbert in a famous communication first came out with his idea of a “Metamathematik” whose object it was to prove that the mathematical science itself was free from contradictions. In this congress there was the first opportunity for us Germans to meet so many mathematicians from other European countries whom we had known before the war (not I myself, who then was only 24 years of age). So when we met in the evenings in the Burgkeller (a famous Leipzig Restaurant) there was much questioning of the type: What about Prof. X from A., is he still alive ? On such an occasion I happened to be seated together with other young people quite near to the table where Hilbert was seated together with the cream of participants. I heard him put exactly the above type of question to a Hungarian mathematician about another Hungarian mathematician. The former began to answer: “Yes, he teaches at B. now, he occupies himself with the theory of ..., he was married a few years ago, there are three children, the oldest ...” But Hilbert already during the first few words of this long tirade began zu say repeatedly: “Ja, aber.., ja aber, ...”, and when he eventually succeeded in stopping the other’s flow of words, he continued: “Ja, aber, das wollte ich ja gar nicht alles wissen, ich habe doch nur gefragt ‘existiert er noch?’ ” And the word “existiert” was most emphatically pronounced. Now as to your request to have a few words on my personal reaction of Hilbert’s work in the theory of numbers, I am afraid, that would be rather difficult for me. My own attitude towards theory of numbers is exactly opposite to that speaking out of that last story. Whereas he was always interested only in questions of pure existence, I am much in favour of all possible sorts of detail. This shall not mean, that I do not understand and even value rather high an attitude like Hilbert’s. But I cannot be very enthusiastic about it. I think I have done the best I could do in that direction in my Foreword to vol.1 of Hilbert’s Gesammelte Abhandlungen, Berlin 1932, where I gave a survey on his papers in the algebraic number theory. Perhaps you look that up first. I think you will find there what you want.

With kind regards, Yours very sincerely,

1.5 10.01.1968, Reid an Hasse

H. now back in Germany. R. has finished the manuscript on Hilbert. Where are Hilbert’s letters to Minkowski? Does H. know about the circumstances of Hilbert’s name appearing on the proclamation in favor of Hitler?

70 Piedmont Street, San Francisco, California, USA 94117

January 10, 1968

Dear Professor Hasse:^{1}

No doubt you are back in Germany now and enjoying the snow as a change from the sand! I have finally finished my manuscript on Hilbert and am now checking and revising it preparatory to making a copy for Professor Bernays and Professor Courant to read. There is still one great lack. Although I have the letters which Minkowski wrote to Hilbert during the course of their friendship, I have never been able to locate the letters which Hilbert wrote to Minkowski. In the course of time, both sets of letters were returned to the families of their respective authors, so the Hilbert letters should be with the Hilbert material at the Institute in Göttingen; but they are not. (Since I last wrote to you, I have been in Göttingen and had an opportunity to examine this material personally.) The closest thing to a clue which I have received was from Arnold Schmidt, whom I saw in Marburg just a few months before his recent death. He told me that during the war, after Hilbert’s death, Mrs. Hilbert placed his papers in the custody of a young man who was not among those who were close to Hilbert. It was Schmidt’s opinion that the letters were destroyed through inadvertence; but it seems strange that a large packet of letters should have been destroyed when even such trivia as Hilbert’s blood count in 1938 has survived. I am wondering if you have any idea who the person Schmidt mentioned could have been? He himself had forgotten the name. I would very much appreciate any information of suggestions you can give. You have been so helpful already that I hate to bother you again; but I have a recurrent nightmare that the day after the book is published those letters will turn up!

1Notiz von Hasse: Grunwald vorlegen

I have one last question. Do you, by any chance, know anything of the circumstances of Hilbert’s name appearing on the enclosed statement? From all that I have gathered from people close to him, signing such a statement would not have been consistent with his character of his views.^{2} May I again express my gratitude for the assistance which you have given me?

Sincerely yours,

Constance Reid

2Reid refers to the proclamation in favor of Hitler ^{▶}, see page 36.

1.6 28.01.1968, Hasse an Reid

H. asked Grunwald about Hilbert’s letters to Minkowski. About the 1934 proclamation.

January 28, 1968

Dear Mrs. Reid:

I am very sorry I cannot help you with your question concerning the bunch of letters Hilbert wrote to Minkowski. I have no idea who the young man might be, to whom Arnold Schmidt gave the letters.

I let Professor Dr. Wilhelm Grunwald—the director of the University Library Göttingen—read your letter. He is a former pupil of mine and we are good friends now. He let one of his people go through the Minkowski ‘Nachlaß’, but the letters in question were not found there. He has further asked one of his colleagues to pursue the trace mentioned by Arnold Schmidt. I do not know what he means by this, but I should think he has a definite idea what could be done. If anything of interest turns out, he will let me know, and I shall forward the information to you.

With a good deal of the names under the 1934 proclamation^{▶} Hilbert is in excellent company. Scientists like Abderhalden, Bier, Esau, Hamel, Hartmann, Heidegger, Jaensch, Kisch, Krüß, Martius, Panzer, Schmeidler, Tammann, Trendelenburg, Valentiner have been patriotic Germans but are above all suspicion of leaning towards national–socialism.

Yours very sincerely,

H. Hasse

1.7 21.01.1970, Reid an Hasse

R. sends H. her book on Hilbert, with many thanks.

70 Piedmont Street, San Francisco, Calif., USA

January 21, 1970

Dear Professor Hasse,^{1}

The book on Hilbert—with which you gave me so much assistance—is being published this week; and I have asked Springer Verlag to send you a copy with my compliments. Since, however, the book will be coming to you directly from Heidelberg, I would like to add my personal thanks for the time and effort you expended for me. You were particularly helpful; and as you will see, I have been able to use most of the things you told me about Hilbert’s last years to very good advantage.

I hope that you enjoy the book and find it an accurate picture of Hilbert during the period when you were close to him.

Sincerely yours,

Constance Reid

1Notiz von Hasse: Warten bis Buch gelesen

1.8 01.02.1970, Hasse an Reid

Thanks for the book. Some comments. About Hilbert’s lecture course 1919 on “Anschauliche Geometry”. H. explains his political attitude. H. had not been a “nationalist” but considered himself as a national citizen. H. had never been a member of any political party.

February 1, 1970

Dear Mrs. Reid,

Thank you for your kind letter^{▶}. It was a great pleasure for me to help you with the Hilbert book.

In the meantime, the book itself has arrived from Heidelberg. Thank you very much indeed for letting me have it. I have already read considerable parts, particularly those about the time when I was a student at Göttingen and those about Hilbert’s last years. I find it quite excellent. You have achieved a wonderful piece of work. The whole mathematical world owes you a profound gratitude for unrolling once more so vividly this mastermind’s life.

There are only very few points I should like to draw your attention to.

i) p.206. – Siegel never has been a professor at Hamburg. Before coming to Göttingen in the late thirties he was a professor at Frankfurt.

ii) p.199, l.9–14. – Even in 1919 when I was a student at Göttingen and attended Hilbert’s course of lectures on “Anschauliche Geometrie”, he very often got “stuck” and cast his helpless eyes towards Bernays, his assistant, who had prepared the lectures for him and attended every single one. Bernays then had to contribute the missing link in the argumentation or unravel an incomplete chain of incoherent arguments.

iii) p.207, l.5. – I somehow resent your statement that I “had long been a convinced nationalist”. This term has nowadays a deteriorating sound, and I do not think it applies to me at any time of my life. My feelings have always been, and still are, those of a truly national citizen. I have never belonged to any political party, but like the great majority of my countrymen I resented deeply the injustice done to us by the Treaty of Versailles. There were plebiscites whereever it would be presumably to our damage, but there were none where we clearly would have won: South Tirol and Alsace. This was against what had been promised to us by your President Wilson and had made us lay down our arms. I was however strongly opposed to all nationalistic excesses, particularly to Hitler’s antisemitic policy. For this reason the students revolted against me when I came to Göttingen in May 1934 on the proposal of Hermann Weyl who then had left for your country.

If your wonderful book will see a second edition, as I hope, nay, am sure, it will, I should be greatly obliged to you for putting these points straight.

Yours very sincerely,

H. Hasse

PS. I hope this letter will reach you although I do not have your ZIP code.

1.9 20.02.1970, Reid an Hasse

R. has already a file of corrections and would appreciate receiving addtional remarks. The word “nationalist” is not meant in a derogatory sense.

70 Piedmont Street, San Francisco, Calif., USA

February 20, 1970

Dear Professor Hasse:

Thank you for your kind letter about the Hilbert book^{▶}. I am very happy that you enjoyed it. In respect to the statement on p.207, l.5, to the effect that you “had long been a convinced nationalist”, I assure you that I meant by this to denote exactly the position which you described in your letter^{▶} as yours. However, in discussing the matter the other day with Dr. Peters of Springer Verlag, I learned also from him that in Germany today the word “nationalist” has a derogatory meaning that it does not have in this country. If I had known this, I would not have used it as I did. I have already established a file of corrections and additions in preparation for the time when I have an opportunity to revise my work and I would appreciate receiving any additional remarks which occur to you in the course of your reading. Thank you again for the assistance you have already given me.

1.10 29.12.1970, Reid an Hasse

H. is in San Diego. R. would like to meet H. and talk further about Hilbert in his last years.

70 Piedmont Street, San Francisco, California 94117

December 29, 1970

Dear Professor Hasse:^{1}

I have recently learned from my sister and brother–in–law, Julia and Raphael Robinson, that you are at San Diego State College this year. It seems a most fortunate opportunity for us to meet and talk further about Hilbert and also about Göttingen during his last years. Will you be in San Diego until June, or only until February? I would prefer to come to San Diego this spring; but if you are leaving by the earlier date, I would of course arrange to come some time this month. I know you will be pleased to hear that the Hilbert book is doing very well. I have just received a letter from Springer telling me that it is being translated into Japanese. As I wrote to you once, I am continuing to collect material whenever I have the opportunity and look forward to a more exact and an expanded new edition. I am also planning a trip to New York in January so I would appreciate hearing from you at your earliest opportunity.

Sincerely yours,

Constance Reid

1Randnotiz von Hasse: 5.1.71 m. Postk. bedankt, Daten meiner Abwesenheit mitgeteilt.

1.11 18.03.1971, Reid an Hasse

Trying to find a date for meeting Hasse. H. proposes to meet on 24 April at Fibonacci conference in San Francisco.

70 Piedmont Street, San Francisco, Calif. 94117

March 18, 1971

Dear Professor Hasse:^{1}

Thank you for your card of January 5. Your schedule seemed so full that I decided to wait to come until after March 14. But now you probably have still more engagements on your calendar. I would, however, like to suggest several possible dates for our meeting to talk over Hilbert and the Göttingen days:

Monday, March 29

Friday, April 12 perhaps not too good because of the Easter holiday.

Monday, April 12

It would be most convenient for me, because I would be flying down in the morning, to see you in the afternoon at one or two o’clock. I find that an interview usually runs approximately two hours. It also occurs to me that I might save myself the trip by seeing you when you are in the Bay Area for the meeting of the Fibonacci group. If you would have the time and inclination to see me some time that weekend, I could easily drive over to St. Mary’s. But please choose the date and the arrangement which is most convenient for you.

Sincerely yours,

Constance Reid

1Randnotiz von Hasse: 24.3.71 – Treffen am 24.4. bei Fibonacci–Tagung in San Francisco vorgeschlagen. – Sonst 2.4. für mich am besten

1.12 28.03.1971, Reid an Hasse

Will meet H. on 24 April in S.F.

70 Piedmont Street, San Francisco, Ca. 94117

March 28, 1971

Dear Professor Hasse:

Thank you very much for your letter of March 24. I think I shall take a chance on seeing you at the meeting of the Fibonacci Association on Saturday, April 24, at the University of San Francisco, although there is a slight possibility that I may have to be out–of–town on that date. In that case I shall arrange to come down to San Diego some time during the month of May. I shall of course let you know ahead of time if I shall not be able to be at USF when you come. Unless you hear from me differently, however, I am planning to meet you on the morning of 24th at nine o’clock, or whenever the meeting is scheduled to begin. I can easily find that out. I do not think you will have any difficulty in recognizing me, since I look quite a bit like my sister, Julia Robinson, whom I believe you already know. Looking forward to our conversation, I remain

Yours sincerely,

Constance Reid

1.13 17.04.1976, Reid an Hasse

R. finished her book on Courant. R. has asked Springer to send galley proofs to H., so he would be able to see how R. treats H.’s role in Göttingen after 1933.

70 Piedmont Street, San Francisco 94117

April 17, 1976

Dear Professor Hasse:^{1}

I have finally finished my book about Courant. It will be published in August or September by Springer Verlag. I know that you are interested in my treatment of your role in Göttingen after 1933, and I have asked the people at Springer to send you galley proofs of the sections in which you are mentioned. These will come to you sometime after the beginning of May. I want to take this opportunity to tell you again how appreciative I am of the time and help you have given me in connection with both of my books about Göttingen.

Sincerely yours,

Constance Reid

cc Springer (2)

1Vermerk von Hasse: Korrekturfahnen abwarten

1.14 09.05.1976, Hasse an Reid

H. thanks for Galley proofs. Looking forward to have full story of Courant’s life. 9.5.1976

Dear Mrs. Reid,

Thank you for letting me have the two galley proofs of your Courant Biography on which you mention me.

I have nothing to remark. Everything is correct, so far as I can still remember, and your representation is lucid and to the point. I am looking forward to have the full story of Courant’s life.

With kind regards and good wishes, Yours sincerely,

1.15 11.05.1976, Hasse an Reid

H. has received more galley proofs. H. points out that there are statements which are not correct and asks for correction. About H.’s application for membership in the Nazi party; this was never granted. About H.’s activities in Göttingen. H. could help mathematicians who had problems of political nature. About the happenings in Göttingen in 1945/46.

May 11, 1976

Dear Mrs. Reid,

My letter to you of May 9^{▶} concerned only galley proofs 93 and 94, to– day I got from Springer also galley proofs 98, 99, 111, 137, 144. On each one of the latter there are statements which are against or far from the truth. I must urgently ask you to correct them.

On sheet 111, second half, you say that by this time I was a member of the Nazi party. This is not true I have never been accepted as a member of that party. The truth is as follows.

My endeavour at that time was to keep up Göttingen’s mathematical glory. For this I needed the consensus of the Party Functionaries, whenever I wanted to get some distinguished mathematician for a free position in Göttingen. Amongst those Functionaries I had one close friend and one who was leaning towards helping me. They asked me to join the party; they could help me then better. It is true that I gave in and applied for membership. But in my application I filled in that there was a yewish branch in my father’s family. I was almost sure that this would lead to my application being declined. And so it was. The answer was, that the application was put back until the war was over. I got this answer only after the outbreak of the war. In the meantime I could help mathematicians like Hans Rohrbach, who had political differences under Bieberbach at Berlin, Martin Eichler, who was in a similar situation in Greifswald Halle, and Carl Ludwig Siegel, who had become politically impossible in Frankfurt, by offering them positions in Göttingen.

Also on sheet 111, second half, there is the remark of what I assured to Artin. I cannot remember that I did this. It may be, but certainly not as an “assurance”. How could I do that. I only knew that such cases existed.

On sheet 137, first half: When I re–entered the Mathematics Institute in September 1945, after having been dismissed from American prisonership of war, the rooms were in no way “busy with allied lawyers, interrogating the population”—I had free access to my rooms there, and also to all the other rooms. One day American or English officials (not in uniform) came and wanted to have some book from our Library. I asked them whether they wanted to borrow it. “No” they said, “we want to take it away for ever”. I protested, but they said we Germans had behaved like that in France, and so it was only fair that they could do the same here.

I was not dismissed from my office in the University at that time. This happened only in the following year 1946 (about February). I do not know up to this day, whether it was the consequence of the conversation with two Americans I had in my home at Göttingen towards the end of 1945, or of my remark in the first meeting of our Faculty, mentioned on sheet 137 shortly after.

As to the conversation with two Americans at my home: They wanted to have the manuscript of my investigations during the war on a ballistic problem. I served in the German Admiralty as scientific officer during the second war, having been a young Marine Officer during the first War. I told them that I had burned my manuscript when the occupation army was near to Göttingen. In the following conversation with those American officers (and not in that first Faculty meeting) I made the remarks of “reeducation of Germany”. In the first Faculty meeting I protested only against making the same mistake as Hitler, in the opposite sense.

Finally, on sheet 144, the remark that I had joined the Party must be removed. I suppose that Siegel has brought it up, as the consequence of the fact that as long as I was an applicant for membership I had to wear the party emblem.

Dear Mrs. Reid, I am afraid you will have to rewrite the passages on sheets 137 and 144. If in any way possible, let me have the new text before it goes to print, please.

Asking you to show understanding for my position, Yours very sincerely, H. Hasse

PS. Please excuse my bad typing, with the great number of corrections.

1.16 12.05.1976, Hasse an Reid

Additional remarks to yesterday’s letter. H. explains his political feelings. He has never been a member of any (political) Party. H. explains his reaction at the Faculty meeting in 1945. H. was strongly opposed to the antisemitism of Hitler and his party. H. names some jewish friends, among them Hensel, Courant and Toeplitz.

May 12, 1976

Dear Mrs. Reid,

Allow me to add a few lines to what I wrote you yesterday^{▶}. I was so perplexed and stirred by what you said about me in your Courant Biography that I could hardly concentrate my thoughts, all the more as I had to explain the real facts in a foreign language.

My political feelings have never been National–Socialistic, but always National in the sense of the Deutschnationale Partei of those times, which succeeded the Conservatives of the Second Empire (under Wilhelm II). But I have never been a member of any Party. Inner politics have never interested me, nor do they now. In outer politics I had strong feelings for Germany as it was created by Bismarck in 1871. When this was heavily damaged by the Treaty of Versailles in 1919, I resented that very much. I very much approved Hitlers endeavours to remove the injustices done to Germany in that Treaty, e.g., the so–called Polish Corridor, and also that the rest of the remaining Austria–Hungary was forbidden, to unite with Germany. When Hitler achieved this latter unification, I approved with all my heart and soul. It was the completion of Bismarck’s work in 1871.

It was from this truly national standpoint, that I reacted in that Faculty Meeting after the second war, when the Faculty more or less suggested that one should not admit such a standpoint for its members. And it was also on the background of my remarks to the two Americans who came to my house.

For each one that really knew me it was absurd to connect me with the Antisemitism of Hitler and its party. I was strongly opposed to it and tried to help where I could, not only because of that yewish great–great–grandmother of mine, but much more because the inhumanity of the whole thing, and because I had quite a few yewish friends and amongst them my adored teacher and Ph.D. father Kurt Hensel, grandson of the sister of Felix–Mendelssohn, and also Richard Courant who was particularly nice to me when I came to Göttingen in 1919 as a young student, and also Otto Toeplitz who gave me my first academic position at Kiel in 1922.

I very much hope, dear Mrs. Reid, that you will correct your text in such a way that the readers will no longer have the impression that I had been also “one of those Nazis”, as people say it so often and carelessly. So much for to–day. Thanks again for giving regard to my corrections and supplements.

Please correct in my yesterday’s letter the name of the university town from which I took Martin Eichler to Göttingen. It was not Greifswald, but Halle.

With kind regards, Yours very sincerely (in the true sense of the word)

1.17 19.05.1976, Reid an Hasse

R. accepts H.s statement regarding his status as party member. R. encloses a statement of the Berlin Document Center on that point. R. has tried to present controversial material—which comes up frequently in connection with Courant himself—from the point of view of both sides. 70 Piedmont

Street, San Francisco 94117

May 19, 1976

Dear Professor Hasse,

I have received your three letters regarding the mention of you in my book about Courant^{▶}^{▶}^{▶}. I am very sorry to have caused you so much distress.

In connection with my book on Hilbert, many people objected to my description of you as “a convinced nationalist.” They demanded, “Why didn’t you just say he was a Nazi?” I replied that you had assured me in a letter (2-1-70)^{▶} that you had never belonged to any political party.

Since I knew that I would be referring to you again in my book about Courant and with the intention of obtaining support for the statement you had made, I wrote to the Office of German Affairs in the U.S. State Department for permission to obtain information as to your party affiliation from the Berlin Document Center. I received the enclosed letter^{1}. On the basis of the information contained in it, I made the statements to which you object in #111 and #144. (Incidentally, Siegel has never made any comments to me in this connection.)

I want to assure you with this letter that I accept your statement regarding your status as a party member and have revised—or, rather, am in the process of revising my manuscript accordingly by quoting directly from you. I wanted to send the changes with this letter, but I have not had time to work on them, since I have been finishing up the galley proofs for the first part of the book. They will come to you shortly.

I hope you can see from the excerpts which you have read that throughout the book I have tried to present controversial material—which comes up frequently in connection with Courant himself—from the point of view of both sides and without any attempt on my part to make a final judgment.

1vielleicht den vom 13.4.1972^{▶}

The general attitude of the book in regard to mathematicians in Germany from 1933–1945 is summed up in Friedrich’s remark on #83.

In parts of the book which you did not receive, since they do not deal with the Nazi period, I mention your relationship with Toeplitz during the war and also with Courant in Göttingen immediately after the war.

I will try to mail my revisions this weekend.

1.18 24.05.1976, Reid an Hasse

R. encloses the revisions in her galleys on the basis of H.s letters. Also another paragraph where H. is mentioned in relation to Artin.

70 Piedmont Street, San Francisco CA 94117

May 24, 1976

Dear Professor Hasse,

I am enclosing the revisions I have made in my galleys on the basis of your letters of May 11^{▶} and 12^{▶}.

I have had to shorten and combine statements from both letters, but I hope that you will feel that the result is a fair statement of your point of view and your actual relationship with the National Socialist Party.

I have also included one other paragraph in which you are mentioned in relation to Artin from a galley you did not receive.

Most sincerely,

Constance Reid

1.19 05.06.1976, Hasse an Reid

H. replies and details certain controversial material in her book.

June 5, 1976

Dear Mrs. Reid,

Thank you for your letters of May 19^{▶} and 24^{▶}. Excuse me, please, for answering only to–day. I have been absent from home for the last week of May, and found a terribly large pile of urgent correspondence at my return.

I sincerely appreciate your attitude of “presenting controversial material from the point of view of both sides without any attempt on your part to make a final judgement”. […] remark, that this endeavour of yours would perhaps be even more convincing, if you replaced “Nazi Party” by “National Socialist Party”. Observe that even your Berlin Mission in their letter to you uses the latter official denotation.

As to that letter^{▶}: It is true that I applied for membership in May 1937. But I never received a formal notice of my acception. I was only told orally by my “Ortsgruppenleiter” that my application had been received under the number 5.619.530, and that the final decision would come later. Only four years later I got a notice, that this decision would be postponed until after the war.

As to your new text on # 111, there is only a small point not quite correct, viz., considering the party functionaries who helped me to fill vacancies in Göttingen by distinguished mathematicians regardless of their political leanings. The two such functionaries I mentioned to you in my letter of May 11^{▶} were:

1. Dr. P. Ziegenbein, former student of Prof. Kaluza, Kiel, who was made my chief assistant by Prof. Vahlen (at that time director in the Ministry of Education), with the instruction to calm down the revolting National–Socialist students at Göttingen who wanted a Party member instead of me as successor of Prof. H. Weyl. Dr. Ziegenbein had a very low Party number and correspondingly a great influence on the Party functionaries in Göttingen’s University and in the Ministry.

2. The chief clerk of the Math. Dept. of the Göttingen University, who at the same time was my “Ortsgruppenleiter” mentioned above.

Consequently you have to change the words “Party functionaries in the Ministry” on # 111. You could just replace “Ministry” by University.

As to # 137^{1}, I am really perplexed by Rellich’s letter to Courant. So far as I remember there was no “allied lawyers” at all in the Math. Institute when I returned there when I was dismissed as a prisoner of war in Sept. 1945. My dismission from the Faculty happened only half a year later, in Febr. 1946, whereas the Faculty meeting you mention before was in Sept. or Oct. 1945.

# 144 I am not aware of having been on “Du” with Artin, neither before he emigrated, nor after his return. You would better leave out that remark although it would have been a great honour for me. True is, that Artin and I were on excellent terms both before and after.

So much for to–day. Thanks again for giving regard to my corrections and supplements,

Yours very sincerely,

H. Hasse

1Zahl undeutlich; vielleicht auch ‘111’.

Kapitel 2 Weiteres Material zu Hasse/Reid

\underline{18.08.1934, Proklamation} 36

2.1 18.08.1934, Proklamation

𝔚𝔞𝔯𝔲𝔪 𝔄𝔡𝔬𝔩𝔣 ℌ𝔦𝔱𝔩𝔢𝔯?

𝔇𝔦𝔢 𝔡𝔢𝔲𝔱𝔰𝔠𝔥𝔢 𝔚𝔦𝔰𝔰𝔢𝔫𝔰𝔠𝔥𝔞𝔣𝔱 𝔥𝔦𝔫𝔱𝔢𝔯 𝔡𝔢𝔪 𝔉𝔲^{··}𝔥𝔯𝔢𝔯

dnb Berlin, 18. August. 34

Eine große Anzahl führender deutscher Wissenschaftler hat folgenden Aufruf erlassen:

Am 19. August steht das deutsche Volk erneut vor einer Entscheidung, die über seine Zukunft bestimmen wird. Durch den Entschluß der Reichsregierung, das Amt das Reichskanzlers und Reichspräsidenten in der Person des Führers Adolf Hitler zu vereinigen, ist eine Sorge behoben, die viele deutschen Männer in den Tagen bedrückt hat, in denen das deutsche Volk bangend am Krankenlager des verewigten Reichspräsidenten und Generalfeldmarschalls gestanden hat.

Wir unterzeichneten Vertreter der deutschen Wissenschaft, die wir auch namens vieler sprechen, die in diesen Tagen weder durch Wort noch Brief für uns erreichbar waren, haben das V e r t r a u e n z u A d o l f H i t l e r a l s S t a a t s f ü h r e r, daß er das deutsche Volk aus seiner Not und Bedrückung herausführen wird. Wir vertrauen auf ihn, daß auch die Wissenschaft unter seiner Führung die Förderung erfahren wird, deren sie in ihrer Gesamtheit bedarf, um die hohe Aufgabe zu erfüllen, die ihr beim Wiederaufbau der Nation zukommt.

Um der Wirkung nach innen wie nach außen willen muß erneut die Einheit und Geschlossenheit des deutschen Volkes und seines Willens zur Freiheit und Ehre durch das Bekenntnis zur Führerschaft Adolf Hitlers zum Ausdruck gebracht werden. Die unterzeichneten Vertreter der deutschen Wissenschaft folgen dem Appell der Reichsregierung, mit dem das deutsche Volk am 19. August zur Entscheidung gerufen wird.

Abderhalden–Halle; von Arnim, Rektor der Technischen Hochschule Berlin; Appel–Berlin; Bieberbach–Berlin; Bier–Berlin; Brackmann, Direktor des Geheimen Staatsarchivs, Berlin; Brandi–Göttingen, Vorsitzender des Verbandes deutscher Historiker; Eitel–Berlin; Deichmann–Berlin; Esau, Rektor der Universität Jena; Eugen Fischer, Rektor der Universität Berlin; Herwart Fischer, Rektor der Universität Würzburg Führer des Reichsverbandes deutscher Hochschulen, Föttinger–Berlin; Gocht–Berlin.

Hamel–Berlin; Nicolai Hartmann–Berlin; Haushofer–München; Heidegger–Freiburg; Hergesell–Berlin; Herzog, Rudolf–Gießen; Heymann–Berlin; Hilbert–Göttingen; His–Berlin (Brombach i. Bad.); Horn–Berlin; Jaensch–Marburg; Jander–Berlin; Kahle–Bonn; Kisch–München; Kaellreuter–München; Kreucker–Berlin; Krückmann–Berlin; Krüger, Rektor der Tierärztlichen Hochschule Berlin; Krüger, F. Greifswald; Krüß, Generaldirektor der Staatsbibliothek Berlin.

Lienau–Danzig; Marcks, Erich–Berlin; Martius–Göttingen; Meinhof–Hamburg; Mußka–Marburg; von Müller–München; Panzer–Heidelberg; Pahr–Leipzig; Petersen–Berlin; Prion–Berlin; Rüdin–München; Erich Seeberg–Berlin; Reinhold Seeberg–Berlin; Sombart–Berlin; Schittenhelm–Kiel; Schmeidler–Breslau; Schmieden–Frankfurt a. Main; Schmidt–Eberswalde; Schmitt, Karl, Berlin; Schroeder, Edward, Göttingen; Schwemann–Aachen; Schword–Hannover; Spamer–Dresden; Spangenberg–München; Stark–Berlin; Stickel–Berlin; Stoeckel–Berlin;

Storm–Berlin; Tammann–Göttingen; Thiersch–Göttingen; Trendelenburg–Berlin; Ministerialdirektor Vahlen, Berlin; Valentiner–Clausthal; Wetzel–Greifswald; Theodor Wiegand, Präsident des Archäologischen Instituts Berlin; Ziesemer–Königsberg Pr.

\underline{13.04.1972, Berlin Docum. Center an Reid} 38

2.2 13.04.1972, Berlin Docum. Center an Reid

MISSION OF THE UNITED STATES OF AMERICA

In reply refer to BDC/254/72/RB/tw

Berlin Document Center, 1 Berlin 37 Wasserkaefersteig 1

April 13, 1972

Dear Miss Reid:

In reply to your letter dated April 7, 1972, we have received a letter from Mr. Geoffrey W. Chapman, Office of German Affairs, giving you permission to receive information concerning Prof. Helmut Hasse. The following information was derived from our files at the Berlin Document Center:

NSDAP entry date : May 1, 1937 NSDAP # : 5.639.530

In December 1939 Professor Hasse filed a petition by way of clemency in order to remain a Nazi Party member despite his partly Jewish descent. In April 1941 a decision on his application was deferred until after his dismissal from the German army. Professor Hasse was serving in the German Kriegsmarine as a Kapitaenleutnant at that time. Since his available Party records do not show any interruptions of his membership status it seems that he was listed as a regular member of the NSDAP until the end of WWII.

We hope this information will be of help to you.

Sincerely yours,

Richard Bauer

Deputy Director Berlin Document Center

\underline{03.05.1976, Springer an Hasse} 39

2.3 03.05.1976, Springer an Hasse

Springer–Verlag New York·Publishers May 3, 1976

Dear Professor Hasse:

Enclosed please find galley pages 83 and 88 which Ms. Reid requested that we send you.

Cordially yours,

Ruth Adams

Production Editor

Enc. cc: Constance Reid

2.4 05.05.1976, Springer an Hasse

Springer–Verlag New York·Publishers

May 5, 1976^{1}

Dear Professor Hasse:

Enclosed please find galley pages 93, 94, 98 and 99 which Ms. Reid requested we send you.

Cordially yours, Ruth Adams Production Editor

Enc. cc: Constance Reid

1Vermerk von Hasse: Für die Fahnen 83,88 hatte ich Mrs. Reid bereits am 9.5. gedankt

2.5 07.05.1976, Springer an Hasse

Springer–Verlag New York·Publishers May 7, 1976

Dear Professor Hasse:

Enclosed please find galleys 137 and 144 which Ms. Reid requested that we send you.

Cordially yours,

Ruth Adams

Production Editor

Enc. cc: Ms. Constance Reid

Kapitel 3 Register

42 Abderhalden, 16 Artin, 26, 32, 34

Bauer, R., 38 Bernays, 6, 14, 18 Bieberbach, 26 Bier, 16 Bismarck, 28

Chapman, 38 Courant, 6, 14, 24, 25, 28–30, 32, 34

Eichler, 26, 29 Esau, 16

Friedrich, 31

Grunwald, 14, 16

Hamel, 16 Hartmann, 16 Hasse, H., 38 Hasses Ur–Urgroßmutter, 28 Hecke, 7 Heidegger, 16 Hensel, 29 Hilbert, 6, 7, 9, 11–14, 16–18, 20–22,

30 Hilbert, K., 8, 14

Jaensch, 16

Kaluza, 33 Kisch, 16 Krüß, 16

Martius, 16

Mendelssohn–Bartholdy, 29 Minkowski, 14, 16

Panzer, 16 Peters, 20

Reid, C., 39–41 Reid, L.W., 6, 7, 11 Rellich, 34 Robinson, J., 21, 23 Robinson, R., 21 Rohrbach, 26

Schmeidler, 16

Schmidt, A., 8, 14, 16 Siegel, 18, 26, 27, 30 Springer, 17, 20, 21, 24

Tammann, 16 Toeplitz, 29, 31 Trendelenburg, 16

Vahlen, 33 Valentiner, 16

Weyl, 19, 33 Wilson, 19

Ziegenbein, 33